# Multilevel Fast Multipole Algorithm for Electromagnetic Scattering of Metasurfaces using a Static Mode Representation

### v1

@@ Abstract
Metasurfaces, comprising large arrays of interacting subwavelength scatterers, pose significant challenges for general-purpose computational methods due to their large electric dimensions and multiscale nature. This paper introduces an efficient boundary element method specifically tailored for metasurfaces, leveraging the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) formulation. Our method combines the Multilevel Fast Multipole Algorithm (MLFMA) with a representation of surface current density unknowns using static modes, a set of entire domain basis functions dependent only on object shape. The reduction in the number of unknowns by the static mode expansion (SME), combined with the $\mathcal{O}(N\log N)$ complexity of MLFMA matrix-vector products, significantly reduces CPU time and memory requirements for solving scattering problems from metasurfaces compared to classical MLFMA with RWG basis functions. The MLFMA-SME method offers substantial benefits for the analysis and optimization of metalenses and metasurfaces.
@@ I Introduction
Metasurfaces are collections of interacting subwavelength scatterers, typically operating near resonance. By carefully designing the geometry and arrangement of these scatterers, metasurfaces can engineer effective constitutive relations. They are also versatile in shaping and manipulating electromagnetic waves for various applications, including imaging and analog computing [1]. Metalenses are one common class of metasurfaces, in which the unit cells are designed to impart a specific, position-dependent phase shift to transmitted light, resulting in a global focusing effect. Accurate simulation of metasurfaces and metalenses, with efficient use of time and memory, is crucial for understanding their fundamental properties and optimizing their performance.
@@
Metasurfaces design often relies on the “unit cell” approximation [2, 3]. This approach constructs a library of the electromagnetic response of meta-atoms, such as disks, pillars, or nanofins, as a function of a design parameter, like the disk radius, pillar height, or the nanofin orientation, assuming the structure to be locally periodic. Unfortunately, the “unit-cell” approach has several limitations: the most critical limitation is the assumption that the interaction between adjacent meta-atoms is similar to that in periodic arrays, which necessitates a slowly changing structure. This constraint limits the design space, resulting in narrow-bandwidth operation, low efficiencies and many types of non-ideality.
@@
General-purpose electromagnetic simulation tools, based on either integral or differential formulation, face significant challenges. Differential formulations are easy to implement and result in sparse matrices but require truncation of the computational domain through the introduction of convenient boundary conditions. Among them, the Finite-Difference Time-Domain (FDTD) is widely used to model metasurfaces. However, the finite resolution leads to inaccuracies in wave speed, resulting in phase accumulation errors as the size of the scattering region increases [4]. Hence, even though the complexity of FDTD simulations nominally increases linearly with the size of the array, maintaining a fixed error requires increasing the resolution. This makes the approach prohibitive when applied using traditional computing platforms.
@@
Integral formulations are appealing because they define the unknowns only within the objects’ volumes or on their boundaries if the objects are spatially piecewise homogeneous, and naturally satisfy radiation conditions at infinity. However, they usually involve dense matrices, which are computationally demanding to invert, even with acceleration techniques like the fast multipole method [5, 6].
A brute-force strategy for scaling full-wave simulations involves adapting general-purpose computational electromagnetic algorithms to massive parallelization for CPU-based [7, 8, 9, 10, 11, 12] or GPU-based computing platforms [13]. However, it is more efficient to tailor existing methods to the specific characteristics of the problem of electromagnetic scattering from metasurfaces, which generally consist of repeated particle shapes arranged in aperiodic positions and arbitrary orientations, typically subwavelength.
@@
In the present work, we propose a fast boundary integral method specifically tailored for large arrays of subwavelength scatterers, leveraging the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) formulation [14, 15, 16, 17]. We combine the Multilevel Fast Multiple Method with the boundary integral equation method employing an expansion of the unknowns of the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) equation in terms of a particular class of entire-domain basis functions, defined on the individual meta-atom, and denoted as static modes. As demonstrated in Ref. [18], static modes drastically reduce the total number of unknowns in scattering problems involving arrays of penetrable particles compared to discretization based on sub-domain basis functions, such as RWG.
@@
This work is organized as follows: In Section II, we recall the PMCHWT formulation and its discretization. In Section III, we briefly review the static mode expansion [18], and in Section IV we combine the the fast multipole method with this expansion. In Section V, we first validate the proposed method by comparing, for arrays of spheres, the results with those obtained using the multiparticle Mie theory [19]. We then assess the computational burden in terms of CPU time and memory requirements by comparing the static mode expansion to a more traditional approach using RWG basis functions. Finally, we present results for three large particle arrays: a golden angle spiral array, a two-layer Moiré superlattice, and a $100\lambda\times 100\lambda$ canonical metalens.
II PMCHWT Surface Integral Equation
@@
Let us consider a system composed of $p$ particles illuminated by a time-harmonic electromagnetic field $\mathrm{Re}\{\mathbf{E}_{\text{inc}}e^{j\omega t}\}$. Each particle is a linear, homogeneous, isotropic material that occupies the three-dimensional domain $\Omega_{\mu}$, with $\mu=1,...,p$. The material has permittivity $\varepsilon^{+}(\omega)$, permeability $\mu^{+}(\omega)$ and it is surrounded by a background medium with permittivity $\varepsilon^{-}(\omega)$ and permeability $\mu^{-}(\omega)$. The equivalent electric $\mathbf{j}_{e}^{\mu}$ and magnetic $\mathbf{j}_{m}^{\mu}$ surface current densities, defined on $\partial\Omega_{\mu}$, are solutions of the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) [14, 15, 16, 17] surface integral problem
@@
where $\delta_{\mu\nu}$ is the Kronecker delta, and
@@
$\mathbf{e}_{0}^{\mu}=-\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\mathbf{E}_{\text{inc}}|_{\partial\Omega_{\mu}},\hskip 17.22217pt\mathbf{h}_{0}^{\mu}=-\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\mathbf{H}_{\text{inc}}|_{\partial\Omega_{\mu}}.$ (4)
The operators $\mathcal{T}_{\pm}^{\mu\nu}$ and $\mathcal{K}_{\pm}^{\mu\nu}$ are, respectively, the EFIE and MFIE integral operators, defined as
@@
$jk^{\pm}\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\int_{\partial\Omega_{\nu}}g^{\pm}(\mathbf{r}-\mathbf{r}^{\prime})\mathbf{u}(\mathbf{r}^{\prime})\,\mathrm{d}S^{\prime}+$
$\frac{1}{jk^{\pm}}\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\int_{\partial\Omega_{\nu}}\nabla g^{\pm}(\mathbf{r}-\mathbf{r}^{\prime})\nabla^{\prime}_{S}\cdot\mathbf{u}(\mathbf{r}^{\prime})\,\mathrm{d}S^{\prime},$ (5)
@@
with $k^{\pm}=\omega\sqrt{\mu^{\pm}\varepsilon^{\pm}}$ and $\zeta^{\pm}=\sqrt{\mu^{\pm}/\varepsilon^{\pm}}$.
Apart from a multiplicative factor, the operator $\mathcal{T}_{\pm}^{\mu\nu}\{\mathbf{u}\}(\mathbf{r})$ takes a surface electric current density $\mathbf{u}$ defined on the $\nu$-th particle and outputs the tangential component of the electric field it generates on the $\mu$-th particle. Similarly, the operator $\mathcal{K}_{\pm}^{\mu\nu}\{\mathbf{u}\}(\mathbf{r})$ takes a surface magnetic current density $\mathbf{u}$ defined on the $\nu$-th particle and outputs the tangential component of the electric field it generates on the $\mu$-th particle.
@@ II.1 Galerkin equations
Aiming at the solution of the PMCWHT eq. (1), we represent the equivalent electric and magnetic surface currents on the particle $\Omega_{\mu}$ as a combination of a set of basis function $\mathcal{B}_{\varphi}^{\mu}=\{\bm{\varphi}^{\mu}_{n}(\mathbf{r})\}_{n=1}^{N}$, so:
$\begin{cases}\displaystyle\mathbf{j}_{e}^{\mu}(\mathbf{r})=\sum_{n=1}^{N}\alpha^{\mu,e}_{n}\bm{\varphi}^{\mu}_{n}(\mathbf{r})\\ \displaystyle\mathbf{j}_{m}^{\mu}(\mathbf{r})=\sum_{n=1}^{N}\alpha^{\mu,m}_{n}\bm{\varphi}_{n}^{\mu}(\mathbf{r}).\end{cases}$ (8)
We find the finite-dimensional approximation of the PMCHWT by substituting eq. (8) in eq. (1) and by projecting along the same set of basis functions, according to a Galerkin projection scheme:
@@
$\mathbf{J}^{\mu}_{e}=[\alpha_{1}^{\mu,e},\ldots,\alpha_{N}^{\mu,e}]^{\mathrm{T}},\hskip 21.52771pt\mathbf{J}_{m}^{\mu}=[\alpha_{1}^{\mu,m},\ldots,\alpha_{N}^{\mu,m}]^{\mathrm{T}},$ (13)
@@
If the particles are identical, each block $Z^{\mu\nu}$ is a square matrix of dimension $2N$. Therefore, the finite-dimensional system (9) has $2Np$ degrees of freedom, making the resolution of problems with a large number of particles prohibitive when using sub-domain basis function as the RWG [20] (where $N$ is the number of edges on the surface mesh of the particle), even when acceleration techniques like fast multipole algorithms are implemented. In the following, we denoted as $\bm{\varphi}$ the sub-domain basis function, such as the Rao-Wilton-Glisson [20].
Let us introduce another set of basis functions $\mathcal{B}_{\psi}^{\mu}=\{\bm{\psi}^{\mu}_{n}(\mathbf{r})\}_{n=1}^{N}$. These functions are related to the basis $\mathcal{B}_{\varphi}^{\mu}$ through the linear transformation
$\bm{\psi}_{i}^{\mu}(\mathbf{r})=\sum_{j=1}^{N}m_{ij}\bm{\varphi}^{\mu}_{j}(\mathbf{r}),$ (16)
@@
where $M=[m_{ij}]$ is the transformation matrix. To find the representations of the operators $T_{\pm}^{\mu\nu}$ and $K_{\pm}^{\mu\nu}$ in the new basis $\mathcal{B}_{\psi}$, we can transform them from the basis $\mathcal{B}_{\varphi}$ representation using
$\{A_{\pm}^{\mu\nu}\}_{\psi}=M\{A_{\pm}^{\mu\nu}\}_{\varphi}M^{\mathrm{T}},$ (17)
where $\{A_{\pm}^{\mu\nu}\}_{\varphi}$ denotes the representation of the operators $A_{\pm}^{\mu\nu}=T_{\pm}^{\mu\nu},K_{\pm}^{\mu\nu}$ in the basis $\mathcal{B}_{\varphi}$.
III Static Surface Current Modes
@@
Guided by the fact that the particles constituting the building blocks of a metasurface, i.e., the meta-atoms, are typically smaller than the wavelength, we seek a set of entire-domain basis functions defined on the surface of an individual meta-atom that can efficiently describe the electric and magnetic surface current densities in this regime. Mathematically, we aim for basis functions that diagonalize some of the operators involved in the PMCHWT formulation in the static limit. This set of entire-domain basis functions, defined on the surface of an individual meta-atom and termed static surface current density modes, has recently been introduced and proven effective in representing surface current density fields [21, 18]. These modes consist of two distinct subsets: longitudinal modes, characterized by a zero surface curl, and transverse modes, with zero surface divergence. These subsets are constructed by introducing two separate auxiliary eigenvalue problems. These two eigenvalue problems are associated with two operators, $\mathcal{T}_{0}^{\perp}$ and $\mathcal{T}_{0}^{\parallel}$, which are obtained from the static-limit of the first and second operators in Eq. (5), respectively.
@@
Eq. (19) can be numerically solved by using as expansion and test functions the star basis functions [23, 24]. Longitudinal static modes constitute a basis for the square-integrable functions defined on $\partial\Omega$ and having zero surface curl, and diagonalize the first operator in Eq. (5) in the small-particle limit.
@@
This equation can be numerically solved by using as expansion and test function the loop functions [23, 24]. Trasverse static modes constitute a basis for the square integrable functions defined on $\partial\Omega$ having zero surface divergence, and diagonalize the second operator in Eq. (5) in the small-particle limit. Several characteristics make the use of static modes appealing: i) The retarded Green’s function, serving as the kernel for integral operators in surface integral formulations, can be decomposed into the static Green’s function, which possesses an integrable singularity, and a distinct term that is a regular function. The integral operators that include the static Green’s function are diagonalized by the static modes, thus regularizing the overall problem. ii) Employing the expansion of static modes, along with suitable rescaling and rearrangement of the unknowns, renders the surface integral formulation resistant to the low-frequency breakdown issue. iii) Static modes are solely dependent on the object’s shape, allowing for their universal application across diverse operating frequencies and material properties of the object. This enables a consistent, frequency-independent representation of any scattering scenario involving objects of a particular shape, surpassing other basis sets, such as characteristic modes [25, 26, 14, 27, 28], which vary with frequency and material properties.
@@
As demonstrated in Ref. [18], the transverse static modes $\{\mathbf{j}_{n}^{\perp}\}_{n=1}^{N^{\perp}}$, associated with the first $N^{\perp}$ eigenvalues $\gamma_{n}^{\perp}$ (sorted in descending order), and the longitudinal static modes $\{\mathbf{j}_{m}^{\|}\}_{m=1}^{N^{\|}}$, associated with the first $N^{\|}$ eigenvalues $\gamma_{m}^{\|}$ (sorted in ascending order), can be used as basis function to represent the equivalent electric and magnetic surface currents on $\Omega_{\mu}$, namely $\mathcal{B}_{\psi}^{\mu}=\{\bm{\psi}^{\mu}_{i}\}_{i=1}^{N^{\perp}+N^{\|}}=\{\mathbf{j}_{1}^{\perp},...,\mathbf{j}_{N^{\perp}}^{\perp},\mathbf{j}_{1}^{\|},...,\mathbf{j}_{N^{\|}}^{\|}\}$. The transformation (17) can be used to express the (9) in the static modes basis starting from the RWG representation. For each couple $(\mu,\nu)$ of particles we have
$\hat{T}^{\mu\nu}=\begin{bmatrix}\hat{T}^{\mu\nu}_{\perp\perp}&\hat{T}^{\mu\nu}_{\perp\|}\\ \hat{T}^{\mu\nu}_{\|\perp}&\hat{T}^{\mu\nu}_{\|\|}\\ \end{bmatrix},\hskip 21.52771pt\hat{K}^{\mu\nu}=\begin{bmatrix}\hat{K}^{\mu\nu}_{\perp\perp}&\hat{K}^{\mu\nu}_{\perp\|}\\ \hat{K}^{\mu\nu}_{\|\perp}&\hat{K}^{\mu\nu}_{\|\|}\\ \end{bmatrix},$ (23)
[…]
$\hat{T}^{\mu\nu}=MT^{\mu\nu}M^{\mathrm{T}},\hskip 21.52771pt\hat{K}^{\mu\nu}=MK^{\mu\nu}M^{\mathrm{T}},$ (25)
where $\alpha,\beta=\perp,\|$ and
$M=\begin{bmatrix}M^{\perp}\\ M^{\|}\end{bmatrix}.$ (26)
@@
As the static surface modes are a very effective description of the surface current density field on objects whose dimensions are small compared to the wavelength, the dimension of the operators $\hat{T}$ and $\hat{K}$ is typically smaller than that of $T$ and $K$, the transformation (25) effectively acts as a compression. Consequently, the system (9) now has $2(N^{\perp}+N^{\|})p$ degrees of freedom, which is typically drastically smaller than $2n_{e}p$ derived from the RWG basis, where $n_{e}$ is the number of edges in the single-particle mesh. In the following, we denoted as $\bm{\psi}$ the basis function in the static modes basis.

### v2

# Multilevel Fast Multipole Algorithm for Electromagnetic Scattering by Large Metasurfaces using Static Mode Representation

@@ Abstract
Metasurfaces, consisting of large arrays of interacting subwavelength scatterers, pose significant challenges for general-purpose computational methods due to their large electric dimensions and multiscale nature. This paper introduces an efficient boundary element method specifically tailored for metasurfaces, leveraging the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) formulation. Our method combines the Multilevel Fast Multipole Algorithm (MLFMA) with a representation of the unknown equivalent surface current density by means of static modes, a set of entire domain basis functions dependent only on object shape but independent of the material and frequency. The compression of the number of unknowns enabled by the Static Mode Representation (SMR), combined with the $\mathcal{O}(N\log N)$ complexity of MLFMA matrix-vector products, significantly reduces CPU time and memory requirements compared to classical MLFMA with RWG basis functions. We demonstrate the accuracy, time, and memory requirements of this method through several test cases including the full-wave simulation of a $100\lambda\times 100\lambda$ canonical metalens. The MLFMA-SMR method offers substantial benefits for the analysis and optimization of metasurfaces and metalenses.
@@ I Introduction
Metasurfaces are collections of interacting subwavelength scatterers, typically operating near resonance [1]. By carefully designing the geometry and arrangement of these scatterers, metasurfaces can engineer a wide class of effective constitutive relations [2]. They are also versatile in shaping and manipulating electromagnetic waves for various applications, including analog computing [3], biosensors [4], and imaging [5]. Metalenses are one common class of metasurfaces, in which the unit cells are designed to impart a specific, position-dependent phase shift to transmitted light, resulting in a global focusing effect [6]. The accurate simulation of metasurfaces and metalenses, with efficient use of time and memory, is crucial for both understanding their fundamental properties [7] and optimizing their performance [8].
@@
Metasurfaces design often relies on the “unit cell” approximation [6, 9, 10]. This approach constructs a library of the electromagnetic response of meta-atoms, such as disks, pillars, or nanofins, as a function of a design parameter, like the disk radius, pillar height, or the nanofin orientation, assuming the structure to be locally periodic. Then, the contributions of the different “unit-cell” are combined by linear superposition. Unfortunately, the “unit-cell” approach has several limitations: the most critical limitation is the assumption that the interaction between adjacent meta-atoms is assumed equal to that in periodic arrays, which necessitates a slowly changing structure. This constraint limits the design space, resulting in narrow-bandwidth operation, low efficiencies, and many types of non-ideality.
@@
General-purpose electromagnetic simulation tools, based on either integral or differential formulations, face significant challenges in the simulation of metasurfaces [7]. Differential formulations, including finite-element methods and finite-difference methods, are easy to implement and result in sparse matrices but require truncation of the computational domain through the introduction of convenient boundary conditions [11]. Among them, the finite-difference time-domain (FDTD) method [12, 13, 14] is widely used to model metasurfaces. However, the finite resolution leads to inaccuracies in wave speed, resulting in phase accumulation errors as the size of the scattering region increases [15]. Hence, even though the complexity of FDTD simulations nominally increases linearly with the size of the array, maintaining a fixed error requires increasing the resolution [16, 13]. This makes the approach prohibitive when applied using traditional computing platforms.
@@
Integral formulations are appealing because they define the unknowns only within the objects’ volumes or on their boundaries if the objects are spatially piecewise homogeneous, and naturally satisfy radiation conditions at infinity. However, they usually involve dense matrices, which are computationally demanding to invert, even with acceleration techniques like the fast multipole method [17, 18, 19], hierarchical matrices [20, 21], or QR decomposition [22].
@@
A brute-force strategy for scaling full-wave simulations involves adapting general-purpose computational electromagnetic algorithms to massive parallelization for CPU-based [23, 24, 25, 26, 27, 28] or GPU-based computing platforms [29, 30]. However, there remains significant potential to refine these methods by leveraging the peculiar characteristics of electromagnetic scattering in metasurfaces, which generally involve repeated particle shapes, typically of subwavelength dimensions, arranged in aperiodic positions and arbitrary orientations. In this work, we pursue this strategy by proposing a fast boundary integral method specifically tailored to this problem. We combine the Multilevel Fast Multiple Method with the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) [31, 32, 33, 34] boundary integral equation employing an expansion of the unknowns in terms of a particular class of entire-domain basis functions, defined on the individual meta-atom, and denoted as static modes. As demonstrated in Ref. [35], static modes drastically reduce the number of unknowns in scattering problems involving arrays of penetrable particles compared to discretization based on sub-domain basis functions, such as RWG.
@@
This manuscript is organized as follows: In Section II, we recall the PMCHWT formulation and its discretization. In Section III, we briefly review the static mode basis [35], and in Section IV, we apply the fast multipole method to accelerate the solution of the PMCHWT where the unknown equivalent surface current densities are expanded in terms of static modes. In Section V, we first validate the proposed method by comparing, for arrays of spheres, the results with those obtained using the multiparticle Mie theory [36]. We then assess the computational burden in terms of CPU time and memory requirements by comparing the static mode representation to a more traditional approach using RWG basis functions. Finally, we present results for three large particle arrays: a golden angle spiral array, a two-layer Moiré superlattice, and a $100\lambda\times 100\lambda$ canonical metalens.
II Surface Integral Equation
@@
Let us consider a system composed of $p$ objects illuminated by a time-harmonic electromagnetic field $\mathrm{Re}\{\mathbf{E}_{\text{inc}}e^{j\omega t}\}$, $\mathrm{Re}\{\mathbf{H}_{\text{inc}}e^{j\omega t}\}$. Each object is made of a linear, homogeneous, isotropic material that occupies the three-dimensional domain $\Omega_{\mu}$, with $\mu=1,...,p$. This material is described by a permittivity $\varepsilon^{+}(\omega)$ and a permeability $\mu^{+}(\omega)$. The objects are surrounded by a background medium with permittivity $\varepsilon^{-}(\omega)$ and permeability $\mu^{-}(\omega)$. The equivalent electric $\mathbf{j}_{e}^{\mu}$ and magnetic $\mathbf{j}_{m}^{\mu}$ surface current densities, defined on the boundary $\partial\Omega_{\mu}$ of the domain $\Omega_{\mu}$, are solutions of the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) [31, 32, 33, 34] surface integral problem
@@
where $\delta_{\mu\nu}$ is the Kronecker delta
@@
$\mathbf{e}_{0}^{\mu}=-\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\mathbf{E}_{\text{inc}}|_{\partial\Omega_{\mu}},\hskip 17.22217pt\mathbf{h}_{0}^{\mu}=-\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\mathbf{H}_{\text{inc}}|_{\partial\Omega_{\mu}},$ (4)
and $\hat{n}_{\mu}$ is the normal to the surface $\partial\Omega_{\mu}$, pointing outward. The operators $\mathcal{T}_{\pm}^{\mu\nu}$ and $\mathcal{K}_{\pm}^{\mu\nu}$ are, respectively, the EFIE and MFIE integral operators, defined as
@@
$\frac{1}{jk^{\pm}}\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\int_{\partial\Omega_{\nu}}\nabla g^{\pm}(\mathbf{r}-\mathbf{r}^{\prime})\nabla^{\prime}_{S}\cdot\mathbf{u}(\mathbf{r}^{\prime})\,\mathrm{d}S^{\prime}+$
$jk^{\pm}\hat{\mathbf{n}}_{\mu}\times\hat{\mathbf{n}}_{\mu}\times\int_{\partial\Omega_{\nu}}g^{\pm}(\mathbf{r}-\mathbf{r}^{\prime})\mathbf{u}(\mathbf{r}^{\prime})\,\mathrm{d}S^{\prime},$ (5)
@@
with $k^{\pm}=\omega\sqrt{\mu^{\pm}\varepsilon^{\pm}}$ and $\zeta^{\pm}=\sqrt{\mu^{\pm}/\varepsilon^{\pm}}$. Apart from a multiplicative factor, the operator $\mathcal{T}_{\pm}^{\mu\nu}\{\mathbf{u}\}(\mathbf{r})$ takes a surface electric current density $\mathbf{u}$ defined on the boundary $\partial\Omega_{\nu}$ of the $\nu$-th particle and outputs the tangential component of the electric field it generates on the surface of the $\mu$-th particle. Similarly, the operator $\mathcal{K}_{\pm}^{\mu\nu}\{\mathbf{u}\}(\mathbf{r})$ takes a surface magnetic current density $\mathbf{u}$ defined on $\partial\Omega_{\nu}$ and outputs the tangential component of the electric field it generates on the surface of the $\mu$-th particle.
@@ II.1 Galerkin equations
Aiming at the solution of the PMCWHT of Eq. (1), we represent the equivalent electric $\left\{\mathbf{j}_{e}^{\mu}\right\}_{\mu=1}^{p}$ and magnetic $\left\{\mathbf{j}_{m}^{\mu}\right\}_{\mu=1}^{p}$ surface current densities on the boundary $\partial\Omega_{\mu}$ of the $\mu$-th particle as a linear combination of basis functions $\mathcal{B}_{\varphi}^{\mu}=\{\bm{\varphi}^{\mu}_{n}(\mathbf{r})\}_{n=1}^{N_{\varphi}}$:
$\begin{cases}\displaystyle\mathbf{j}_{e}^{\mu}(\mathbf{r})=\sum_{n=1}^{N_{\varphi}}\alpha^{\mu,e}_{n}\bm{\varphi}^{\mu}_{n}(\mathbf{r})\\ \displaystyle\mathbf{j}_{m}^{\mu}(\mathbf{r})=\sum_{n=1}^{N_{\varphi}}\alpha^{\mu,m}_{n}\bm{\varphi}_{n}^{\mu}(\mathbf{r}).\end{cases}$ (8)
We find the finite-dimensional approximation of the PMCHWT by substituting (8) in (1) and by projecting along the same set of basis functions, according to the Galerkin projection scheme:
@@
$\mathbf{J}^{\mu}_{e}=[\alpha_{1}^{\mu,e},\ldots,\alpha_{N}^{\mu,e}]^{\intercal},\hskip 21.52771pt\mathbf{J}_{m}^{\mu}=[\alpha_{1}^{\mu,m},\ldots,\alpha_{N}^{\mu,m}]^{\intercal},$ (13)
@@
If the particles have the same shape and are discretized in the same way, each block $Z^{\mu\nu}$ is a square matrix of dimension $2N_{\varphi}$. In this case, the finite-dimensional system (9) has $2N_{\varphi}p$ degrees of freedom.
We can choose as basis either sub-domain or entire-domain basis functions. Sub-domain basis functions are non-zero only over a localized region of each individual object. A typical choice for these functions is the Rao-Wilton-Glisson (RWG) functions [37]. We denote the set of sub-domain of basis functions as $\mathcal{B}_{\xi}^{\mu}=\{\bm{\xi}^{\mu}_{n}(\mathbf{r})\}_{n=1}^{N_{\xi}}$. Instead, entire domain basis functions extend over the entire domain of each individual object. These basis functions can be generated through methods like separation of variables (e.g., vector spherical or spheroidal basis functions) or by solving auxiliary eigenvalue problems, as in the case of characteristic modes [38, 31, 39, 40], (see [41, 42]. For further examples, see [41, 42] for arrays of perfectly conducting particles, [43] for perfectly conducting metasurfaces, and material-independent modes [44, 45]. We denote the set of entire-domain basis functions $\mathcal{B}_{\psi}^{\mu}=\{\bm{\psi}^{\mu}_{n}(\mathbf{r})\}_{n=1}^{N_{\psi}}$.
When we use sub-domain basis functions the numerical solution quickly becomes prohibitive as the number of particles $p$ increases, even with the implementation of acceleration techniques such as fast multipole algorithms. This is due to both computational and memory burdens. To address this challenge, we turn to entire domain basis functions. These functions are related to the basis $\mathcal{B}_{\xi}^{\mu}$ through the linear transformation
$\bm{\psi}_{i}^{\mu}(\mathbf{r})=\sum_{j=1}^{N_{\xi}}m_{ij}\bm{\xi}^{\mu}_{j}(\mathbf{r}),$ (16)
@@
where $M=[m_{ij}]$ is a real-valued transformation matrix. To find the representations of the operators $T_{\pm}^{\mu\nu}$ and $K_{\pm}^{\mu\nu}$ in the new basis $\mathcal{B}_{\psi}$, we can transform them from the basis $\mathcal{B}_{\xi}$ representation using
$\{A_{\pm}^{\mu\nu}\}_{\psi}=M\{A_{\pm}^{\mu\nu}\}_{\xi}M^{\intercal},$ (17)
with $A\in\{T,K\}$
III Static Current Density Modes as Entire domain basis
Guided by the fact that the particles constituting the metasurface, i.e., the meta-atoms, are typically smaller than the wavelength, we seek a set of entire-domain basis functions of an individual meta-atom that can efficiently describe the unknown equivalent electric and magnetic surface current densities in the small particle regime. Mathematically speaking, we aim for basis functions that diagonalize some of the operators involved in the PMCHWT formulation in the static limit. We also require that set be complete and thus capable of describing any square-integrable vector function defined on the surface of the object.
@@ III Static Current Density Modes as Entire domain basis
A set of entire-domain basis functions, defined on the surface of an individual meta-atom and termed static surface current density modes, has recently been introduced and proven effective in representing surface current density fields [46, 35]. These modes consist of two distinct subsets: longitudinal modes, characterized by a zero surface curl, and transverse modes, with zero surface divergence. These subsets are constructed by introducing two separate auxiliary eigenvalue problems. These two eigenvalue problems are associated with two operators, $\mathcal{T}_{0}^{\perp}$ and $\mathcal{T}_{0}^{\parallel}$, which are obtained from the static-limit of the first and second operators in the expression of $\mathcal{T}$ (5), respectively.
@@
Apart from a multiplicative factor, the operator $\mathcal{T}_{0}^{\|}$ takes a surface electric current density $\mathbf{w}$ defined on the boundary $\partial\Omega$ of the particle and outputs the tangential component of the electric field it generates on $\partial\Omega$ in the electrostatic limit.
Eq. (18) is numerically solved by using as expansion and test functions the star-basis functions [48, 49, 50]. Longitudinal static modes constitute a basis for the square-integrable functions defined on $\partial\Omega$ and having zero surface curl [51], and diagonalize the first operator in Eq. (5) in the small-particle limit. These modes are the natural modes of electromagnetic scattering from surfaces of finite conductivity in the electrostatic limit [46].
@@
Apart from a multiplicative factor, the operator $\mathcal{T}_{0}^{\perp}$ takes a surface electric current density $\mathbf{w}$ defined on $\partial\Omega$ and outputs the tangential component of the electric field it generates on $\partial\Omega$ in the magnetostatic limit. This equation is numerically solved by using as expansion and test function the loop functions [48, 49, 50]. Transverse static modes constitute a basis for the square integrable functions defined on $\partial\Omega$ having zero surface divergence (can be proved by following Ref. [52]), and diagonalize the second operator in the expression of $\mathcal{T}$ (5) in the small-particle limit. These modes are the natural modes of electromagnetic scattering from surfaces of finite conductivity in the magnetostatic limit [46] .
@@
Several characteristics make the use of static modes appealing. First, the retarded Green’s function, serving as the kernel for integral operators in surface integral formulation, can be decomposed into the static Green’s function, which includes the integrable singularity, and a distinct term that is a regular function. The integral operators that include the static Green’s function are diagonalized by the static modes, thus regularizing the overall problem. Second, employing the expansion of static modes, along with suitable rescaling and rearrangement of the unknowns, renders the surface integral formulation resistant to the low-frequency breakdown issue. Third, static modes are solely dependent on the object’s shape, allowing for their universal application across diverse operating frequencies and materials, under the assumption that the object is homogeneous. This enables a consistent, frequency-independent representation of any scattering scenario involving objects of a particular shape, surpassing other basis sets, such as characteristic modes [41, 38, 31, 39, 40], which vary with frequency and material properties. Fourth, for objects of size smaller than the wavelength of operation, only a few static modes are sufficient to accurately describe the emergent scattering response [35].
@@
The transverse static modes $\{\mathbf{j}_{n}^{\perp}\}_{n=1}^{N^{\perp}}$, associated with the first $N^{\perp}$ eigenvalues $\gamma_{n}^{\perp}$ (sorted in descending order), and the longitudinal static modes $\{\mathbf{j}_{m}^{\|}\}_{m=1}^{N^{\|}}$, associated with the first $N^{\|}$ eigenvalues $\gamma_{m}^{\|}$ (sorted in ascending order) are used as entire domain basis function to represent the equivalent electric and magnetic surface currents on $\Omega_{\mu}$, namely $\mathcal{B}_{\psi}^{\mu}=\{\bm{\psi}^{\mu}_{i}\}_{i=1}^{N^{\perp}+N^{\|}}=\{\mathbf{j}_{1}^{\perp},...,\mathbf{j}_{N^{\perp}}^{\perp},\mathbf{j}_{1}^{\|},...,\mathbf{j}_{N^{\|}}^{\|}\}$.
The matrices of Eq. (9) are first evaluated in terms of the RWG representation. Then the transformation (17) can be used to pass from a RWG representation to the static modes representation.
@@
As the static surface modes are a very effective description of the surface current density field on objects whose dimensions are small compared to the wavelength, the dimension of the operators $\{T\}_{\psi}$ and $\{K\}_{\psi}$ is typically smaller than $\{T\}_{\xi}$ and $\{K\}_{\xi}$, the transformation (17) effectively acts as a compression. Consequently, the system (9) now has $2(N^{\perp}+N^{\|})p$ degrees of freedom, which is typically drastically smaller than $2n_{e}p$ derived from the RWG basis, where $n_{e}$ is the number of edges in the single-particle mesh. In the following, we denote as $\bm{\psi}$ the basis function in the static modes basis.